# Smart Data Portfolios:

### v1

@@ Abstract
Growing concerns about fairness, privacy, robustness, and transparency have made it a central expectation of AI governance that automated decisions be explainable by institutions and intelligible to affected parties. We introduce the Smart Data Portfolio (SDP) framework, which treats data categories as productive but risk-bearing assets, formalizing input governance as an information-risk trade-off. Within this framework, we define two portfolio-level quantities, Informational Return and Governance-Adjusted Risk, whose interaction characterizes data mixtures and generates a Governance-Efficient Frontier. Regulators shape this frontier through risk caps, admissible categories, and weight bands that translate fairness, privacy, robustness, and provenance requirements into measurable constraints on data allocation while preserving model flexibility. A telecommunications illustration shows how different AI services require distinct portfolios within a common governance structure. The framework offers a familiar portfolio logic as an input-level explanation layer suited to the large-scale deployment of AI systems.
Keywords: AI governance, input governance, training data, portfolio theory, explainability, fairness, privacy
@@ 1 Introduction
Artificial intelligence systems increasingly influence access to credit, connectivity, health, education, and public services. When a credit application is denied, or when a child is exposed to harmful recommendations, affected individuals do not ask how gradients propagated through a model. They ask which data were used, in what priority, and under whose governance. While real disputes arise at the level of inputs, existing explainability methods focus on model internals (Strumbelj2014; Lundberg2017; Koh2017), a mismatch that persists even in contemporary AI governance frameworks (Cordeiro2026). This leaves institutions without a principled way to justify which data categories were admissible, how they were used, and how those choices shaped downstream outcomes.
Current regulatory regimes reflect this tension. GDPR Article 15 requires meaningful explanations of automated decisions, the NIST AI Risk Management Framework emphasizes data and input governance (NIST2023), and Article 10 of the EU Artificial Intelligence Act mandates rigorous documentation of training, validation, and testing data (EUAIAct2024). Yet none of these instruments provide procedures for selecting, weighting, or explaining the data that feed high-stakes AI systems. The result is a structural gap between legal expectations of explainability and the technical practice of assembling training corpora. Recent system-level transparency assessments show that even state-of-the-art foundation models provide only limited disclosure of input selection and governance practices (Bommasani2025). The gap widens as AI systems become more adaptive, tool-integrated, and continuously updated in deployment.
A growing body of research points to the same underlying problem. Studies on fairness show that disparate impacts often arise from imbalanced or historically patterned data (Barocas2016; Mehrabi2021). Works on data privacy demonstrate that disclosure risk grows with granularity, linkability, and weak provenance (Narayanan2008; Rocher2019). Robustness failures similarly trace back to narrow or weakly curated training distributions (Recht2019; Taori2020). Taken together, these studies show that high-stakes AI systems require a structured approach to governing the data that underpin model behavior (Sambasivan2021).
@@
We introduce the Smart Data Portfolio (SDP) framework, which treats data categories as productive but risk-bearing information assets whose allocation must be governed, explained, and audited. SDPs provide a dual contribution. First, they transform the qualitative process of data selection into a structured governance problem by defining two measurable quantities: Informational Return and Governance-Adjusted Risk, whose interplay generates a Governance-Efficient Frontier. Second, they supply a model-agnostic explanation layer where institutions justify decisions not by revealing internal model mechanics but by disclosing the regulated portfolio of authorized data. These are documented through Data Portfolio Statements, Data Portfolio Cards, and Consumer Portfolio Reports that record validated portfolio metrics.
This portfolio-based explanation protocol aligns with long-standing practices in financial supervision, where institutions optimize within risk envelopes defined by eligibility rules, concentration limits, and capital buffers. SDPs adopt the same supervisory logic for training data. Regulators set admissible categories, weight bands, and policy risk caps, while institutions select portfolios that maximize informational value subject to these constraints. Crucially, this supervisory structure fits naturally within existing data-governance, risk-management, and compliance processes.
@@
Thus we make four contributions. (1) We develop Smart Data Portfolios as an explanation layer for AI systems, enabling decisions to be justified through governed input portfolios rather than model internals. (2) We formalise informational return, governance-adjusted risk, the governance-efficient frontier, and the policy risk cap, creating a quantitative foundation for input governance. (3) We propose a governance architecture with regulatory tools and reporting artefacts that support transparent and auditable deployment. (4) We illustrate the framework in the telecommunications sector, where heterogeneous applications require distinct governed portfolios under a shared regulatory structure.
@@
The remainder of the paper is organized as follows. Section 2 reviews related research, while Section 3 defines the SDP framework and its principles. Section 4 develops the quantitative framework and derives the governance-efficient frontier. Section 5 introduces the reporting stack that operationalizes SDPs as an explanation layer. Section 6 presents the sectoral illustration, and Section 7 concludes with implications for AI governance.
@@ 2 Background and Related Literature
Research on AI governance spans a wide range of technical, legal, and institutional concerns that have largely evolved in separate strands. Technical literatures have examined fairness, privacy, robustness, and micro-level explainability largely as distinct problems, while parallel work has studied informational value and efficiency. Policy frameworks have only more recently attempted to integrate these perspectives into coherent governance principles. What remains insufficiently developed is a unifying governance approach that can deliver responsible and explainable AI in a form that is consistent, scalable, and intelligible across applications and sectors. This section reviews five strands of research that motivate our focus on input data as a governed portfolio of information assets: (1) transparency and explainability, (2) fairness and bias, (3) privacy and data protection, (4) robustness and data quality, and (5) the value of information and efficiency.
@@ 2.1 Transparency and explainability
An early response to the opacity of machine learning systems focused on explainability at the level of individual predictions. Strumbelj2014 decompose model outputs into feature-level contributions. Lundberg2017 provide a unified framework for explaining outputs based on additive feature attributions, while Koh2017 develop influence functions that trace model behavior back to particular portions of the training set.
Technical approaches to documentation and reporting have complemented these explainability efforts. Mitchell2019 introduce Model Cards as structured summaries of model behavior, including information about training data, evaluation conditions, and subgroup performance. In parallel, dataset documentation frameworks such as Datasheets for Datasets (Gebru2021) record provenance, composition, collection protocols, and intended uses; Paullada2021 survey this literature and highlight its role in improving transparency at the level of individual datasets. Taken together, these contributions provide developers, users, and regulators with clearer insight into how models behave and how particular datasets or examples influence outputs. They also emphasize the value of well-documented datasets, while leaving open the question of which data sources were permissible and under which constraints.
@@ 2.2 Fairness and bias
A large literature examines how machine learning systems can produce unequal or discriminatory outcomes. Barocas2016 analyze how data collection, feature construction, and labeling choices can embed historical inequities into ostensibly neutral models, leading to disparate impacts even in the absence of explicit use of protected attributes. Hardt2016 formalize a notion of equality of opportunity in supervised learning and show that standard training objectives can generate systematic differences in error rates across groups, which motivates explicit fairness constraints and post-processing adjustments. Mehrabi2021 provide a comprehensive survey of bias and fairness in machine learning, cataloging sources of harm that arise from coverage gaps, skewed labels, feedback effects, and mis-specified objectives.
Sectoral studies reinforce these concerns as well. In health care, for example, Chen2021 review ethical machine learning practices and document how imbalanced or unrepresentative training data can translate into clinically significant disparities in performance across patient groups.
These contributions converge on the view that fairness outcomes are tightly linked to the composition and curation of training data, not only to the choice of prediction algorithms. They also suggest that fairness metrics and constraints are most meaningfully evaluated at the level of data mixtures and subgroup coverage than solely at the level of model outputs.
@@ 2.3 Privacy and data protection
Privacy and data protection have long been central concerns in data-intensive systems. Early work on microdata anonymization showed that combinations of quasi-identifiers can sharply increase the risk of reidentification even when direct identifiers are removed (Samarati2001). Subsequent analyses of large-scale data practices demonstrate that linkage across heterogeneous sources creates cumulative disclosure risk, often in ways that are difficult to anticipate (Narayanan2008; Rocher2019). These findings underscore that privacy loss is primarily a function of data structure, granularity, and cross-system interoperability rather than of model design alone.
A complementary line of work develops formal guarantees for limiting individual information leakage. Differential privacy provides mechanisms that bound the inference risk associated with queries or model training and quantifies how privacy loss compounds under composition (Dwork2014; Abowd2018; Wood2020). This literature makes explicit that privacy exposure depends on the magnitude, sensitivity, and reuse of the underlying data, and that even well-designed models can violate privacy when trained on highly identifying inputs.
Together, these strands emphasize that privacy risk is not simply an output-level property of models. It arises from decisions about which data to collect, how long to retain them, how to link them across systems, and how to share or release them. They also imply that privacy requirements will often need to be expressed as constraints on the use and weighting of particular data categories, especially those that are highly identifying or sensitive.
Together, these strands emphasize that privacy risk is not simply an output-level property of models. It arises from decisions about which data to collect, how long to retain them, how to link them across systems, and how to share or release them. They further underscore that governing inferences and data combinations is central to data protection in contemporary AI systems (Wachter2019). As a result, privacy requirements will often need to be expressed as constraints on the use and weighting of particular data categories, especially those that are highly identifying or sensitive.
@@ 2.4 Robustness and data quality
Robustness to changes in data distributions has been the focus of another active line of work. Recht2019 construct a test set for ImageNet and show that image classifiers that perform well on the original benchmark can suffer notable accuracy declines when evaluated on their new but closely related dataset. Taori2020 extend this analysis to a broader range of natural distribution shifts and document systematic performance degradation across models and training regimes. These results suggest that models can be fragile to relatively modest changes in the data generating process, especially when they have been optimized on a narrow distribution of training examples.
In parallel, research on data quality has mapped out common sources of degradation in machine learning procedures. Jakubik2024 review the state of data quality for machine learning and identify issues such as label noise, incomplete or biased coverage, missing documentation, and the inclusion of manipulated or low-veracity sources. They argue that many observed failures of deployed systems can be traced back to shortcomings in data collection and curation rather than to model architecture. This line of work reinforces the idea that robustness and reliability are properties of the combined data and model system, particularly in settings where deployment conditions evolve over time.
@@ 2.5 Value of information and efficiency
A further strand of research considers the value and efficient use of information. Early work in decision analysis, notably Howard1966, develops information value theory and shows that the expected benefit of additional information can be quantified in terms of improved decision quality. Hubbard2014 translates these ideas into practical frameworks for measuring the value of intangibles in organizational settings, emphasizing that many institutions collect more information than they can effectively use.
In finance, Markowitz1952 introduce the idea that portfolio choice can be framed as a balance between expected return and risk, establishing diversification and risk constraints as central features of decision under uncertainty. Subsequent research develops alternative risk measures that capture tail exposures more effectively, most notably Conditional Value-at-Risk (Rockafellar2000). Related work on coherent risk measures (Artzner1999) provides axiomatic foundations for evaluating optimal portfolio allocation. Elton2014 survey the broader development of modern portfolio theory where regulated constraints and model-based estimates jointly determine feasible allocation sets.
In the context of data, Ghorbani2019 assign values to individual data points or subsets based on their marginal contribution to model performance, building on the Shapley value concept from cooperative game theory (Shapley1953). Subsequent work extend these ideas to federated learning, data marketplaces, and scalable privacy-aware settings under heterogeneous data (Wang2020; Sun2023; Tian2022; Wang2023). This line of research shows that not all data are equally valuable, and formalizes information as an asset with quantifiable marginal contributions.
@@ 2.6 Governance frameworks and emerging principles
Policy frameworks have begun to synthesize these research strands into articulated governance aims. The OECD Principles on Artificial Intelligence (OECD2019), the NIST AI Risk Management Framework (NIST2023), the EU Artificial Intelligence Act (EUAIAct2024), and the U.S. Blueprint for an AI Bill of Rights (OSTP2022) all emphasize the need for transparency and documentation of AI systems, fairness and non-discrimination in outcomes, privacy and data protection, robustness and security, and proportional use of data and computational resources. They require, to varying degrees, that high-risk AI systems document training data and model behavior, assess subgroup performance, safeguard privacy, and demonstrate robustness under realistic operating conditions.
Despite these advances, existing governance frameworks largely stop at articulating normative objectives, rather than providing a general architecture capable of sustaining them. They state what AI systems should achieve, but leave unresolved how institutions are to organize, constrain, and explain the data choices that underpin automated decisions. As a result, training data are typically assembled through inherited workflows and ad hoc judgments about value and risk, with limited capacity for consistent external scrutiny across deployments. This need for explanation and justification through input governance motivates the Smart Data Portfolio framework developed in this study.
@@ 3 The Smart Data Portfolio Framework
Building on the preceding literature, the Smart Data Portfolio framework formalizes input governance as a structured allocation problem over regulated data categories. We define a data category as a standardized class of inputs whose internal composition may change over time, but whose use is governed, constrained, and reported as a single entity. Categories are specified by regulators at an intermediate level of detail, capturing shared provenance, governance-risk characteristics, and semantic scope, rather than raw records or model-specific features. This level of abstraction enables data inputs to be governed in a way that is stable, comparable, and auditable across sectors and institutions.
@@
SDPs with different allocations of data categories yield different combinations of informational value and governance-adjusted risk. As a result, some portfolios are strictly superior, in that no alternative provides higher informational return at the same level of governance-adjusted risk. The set of such portfolios for each level of risk defines the Governance-Efficient Frontier. Institutions typically select portfolios on this frontier to maximize informational return subject to the governance constraints, while regulators control the aggregate risk level.
@@
Figure 1: Illustrative governance-efficient frontier under a policy-defined risk cap. Filled markers denote admissible SDPs. Hollow markers denote feasible but dominated SDPs that lie below the frontier. Points to the right of the Risk Cap are policy-infeasible. The governance-optimal portfolio is the highest-return point on the frontier within the admissible region.
@@
These ideas draw on long-established practices in financial supervision. Rather than prescribing valuation models, regulators define admissible regions through eligibility rules, concentration limits, capital buffers, and exposure constraints, while institutions optimize within these boundaries using their own internal methods. The SDP framework adopts this supervisory logic for AI inputs, formalizing how governance parameters shape the feasible design space without constraining model choice or innovation.
4 Methodology
This section formalizes the SDP framework by defining portfolio-level performance and governance risk measures, the resulting governance-efficient frontier, and the regulatory constraints that shape admissible data allocations.

### v2

# arXiv:2512.16452v2[cs.CY] 27 Feb 2026

@@ Abstract
Contemporary AI regulation, including the EU Artificial Intelligence Act and related governance frameworks, increasingly requires institutions to justify the training data used in automated decision-making. Yet existing governance regimes provide limited operational methods for selecting, weighting, and explaining data inputs. We introduce the Smart Data Portfolio (SDP) framework, which treats data categories as productive but risk-bearing assets, formalizing input governance as an information-risk trade-off. Within this framework, we define two portfolio-level quantities, Informational Return and Governance-Adjusted Risk, whose interaction characterizes attainable data mixtures and yields a Governance-Efficient Frontier. Regulators shape this frontier through risk caps, admissible categories, and weight bands that translate fairness, privacy, robustness, and provenance requirements into measurable constraints on data allocation while preserving model flexibility. A sectoral illustration shows how different AI services require distinct portfolios within a common governance structure. The framework provides an input-level explanation layer through which institutions can justify governed data use in large-scale AI deployment.
Keywords: AI governance, input governance, training data, portfolio theory, explainability, privacy, fairness
@@ 1 Introduction
Artificial intelligence systems increasingly mediate access to credit, connectivity, health care, education, and public services. As automated decision-making expands, regulatory attention has shifted from model performance alone toward the governance of the data used to train and operate AI systems. Contemporary regulatory frameworks, including the European Union’s Artificial Intelligence Act (European Parliament and Council of the European Union, 2024) and the NIST Artificial Intelligence Risk Management Framework (National Institute of Standards and Technology, 2023), now increasingly require institutions to document, assess, and justify the quality, representativeness, and provenance of training data. In practice, however, institutions confronting these obligations lack systematic methods for determining how heterogeneous data sources should be selected, combined, and justified within deployed AI systems.
This gap reflects a mismatch between regulatory expectations and prevailing technical approaches to explainability. Individuals affected by automated decisions rarely seek explanations of internal model mechanics. Instead, they ask which data were used, in what priority, and under whose governance. Existing explainability methods, however, focus primarily on model internals or feature-level attribution (Štrumbelj and Kononenko, 2014; Lundberg and Lee, 2017; Koh and Liang, 2017). Institutions therefore remain poorly equipped to justify admissible data choices or to translate governance requirements into operational data-selection decisions, as reflected in persistent transparency gaps around training-data documentation and disclosure practices (Wan et al., 2025).
A growing empirical literature shows that many failures attributed to algorithms originate in the composition and governance of training data. Disparate impacts frequently arise from historically patterned datasets (Barocas and Selbst, 2016; Mehrabi et al., 2021). Privacy risks increase with data granularity and cross-source linkage (Narayanan and Shmatikov, 2008; Rocher et al., 2019). Robustness failures often reflect narrow or weakly curated training distributions (Recht et al., 2019; Taori et al., 2020). Organizational studies further demonstrate how insufficient attention to data practices propagates downstream harms throughout AI development and deployment pipelines (Sambasivan et al., 2021). Taken together, these findings indicate that responsible AI deployment requires structured governance not only of models, but also of the data mixtures that shape model behavior.
@@
To address this problem, we introduce the Smart Data Portfolio (SDP) framework, which treats data categories as productive but risk-bearing information assets whose allocation can be governed, audited, and explained. SDPs formalize input governance as an information-risk trade-off defined at the portfolio level. The framework introduces two measurable quantities, Informational Return and Governance-Adjusted Risk, whose interaction generates a Governance-Efficient Frontier describing admissible data mixtures under regulatory constraints. Rather than explaining decisions through internal model mechanics, institutions justify data use by disclosing governed portfolios of authorized inputs. These portfolios are operationalized through Data Portfolio Statements, Data Portfolio Cards, and Consumer Portfolio Reports.
The SDP framework draws on supervisory principles long established in financial regulation, where institutions operate within regulator-defined risk envelopes while retaining discretion over internal models and strategies. Analogously, regulators specify admissible data categories, governance weight bands, and policy risk caps, while institutions optimize data allocations within these boundaries. This separation between regulatory constraint-setting and institutional optimization provides a model-agnostic mechanism for translating governance requirements into operational data practices compatible with existing compliance and risk-management workflows.
@@
The paper makes four contributions. (1) We develop Smart Data Portfolios as an explanation layer for AI systems, enabling decisions to be justified through governed input portfolios rather than model internals. (2) We formalize informational return, governance-adjusted risk, the governance-efficient frontier, and the policy risk cap, establishing a quantitative foundation for input governance. (3) We propose a governance architecture supported by regulatory tools and reporting artifacts that enable transparent and auditable deployment. (4) We illustrate the framework in the telecommunications sector, where heterogeneous applications require distinct governed portfolios within a shared regulatory structure.
@@
The remainder of the paper is organized as follows. Section 2 reviews related research. Section 3 defines the SDP framework and its principles. Section 4 develops the quantitative framework and derives the governance-efficient frontier. Section 5 introduces the reporting stack that operationalizes SDPs as an explanation layer. Section 6 presents the sectoral illustration. Section 7 concludes with implications for AI governance.
@@ 2 Background and Related Literature
Research relevant to AI governance has developed across technical, economic, and policy domains, often addressing fairness, privacy, robustness, explainability, and informational efficiency as largely separate problems. These literatures have generated substantial insights into the risks and performance characteristics of machine learning systems, yet they typically evaluate models or datasets in isolation rather than the governed allocation of heterogeneous data inputs. As contemporary regulatory frameworks increasingly require institutions to justify how training data are selected and used, the absence of an integrated approach to managing data composition has become more apparent. This section reviews strands of research that collectively motivate treating training data as governed portfolios of information assets.
@@ 2.1 Transparency and explainability
An early response to the opacity of machine learning systems focused on explainability at the level of individual predictions. Štrumbelj and Kononenko (2014) decompose model outputs into feature-level contributions. Lundberg and Lee (2017) provide a unified framework for explaining outputs based on additive feature attributions, while Koh and Liang (2017) develop influence functions that trace model behavior back to particular portions of the training set.
Technical approaches to documentation and reporting have complemented these explainability efforts. Mitchell et al. (2019) introduce Model Cards as structured summaries of model behavior, including information about training data, evaluation conditions, and subgroup performance. In parallel, dataset documentation frameworks such as Datasheets for Datasets (Gebru et al., 2021) record provenance, composition, collection protocols, and intended uses; Paullada et al. (2021) survey this literature and highlight its role in improving transparency at the level of individual datasets. Taken together, these contributions improve transparency regarding model behavior and individual datasets, while leaving largely unresolved how multiple data sources should be selected, combined, and governed within institutional AI deployments.
@@ 2.2 Fairness and bias
A large literature examines how machine learning systems can produce unequal or discriminatory outcomes. Barocas and Selbst (2016) analyze how data collection, feature construction, and labeling choices can embed historical inequities into ostensibly neutral models, leading to disparate impacts even in the absence of explicit use of protected attributes. Hardt et al. (2016) formalize a notion of equality of opportunity in supervised learning and show that standard training objectives can generate systematic differences in error rates across groups, which motivates explicit fairness constraints and post-processing adjustments. Mehrabi et al. (2021) provide a comprehensive survey of bias and fairness in machine learning, cataloging sources of harm that arise from coverage gaps, skewed labels, feedback effects, and mis-specified objectives.
Sectoral studies reinforce these concerns as well. In health care, for example, Chen et al. (2021) review ethical machine learning practices and document how imbalanced or unrepresentative training data can translate into clinically significant disparities in performance across patient groups.
These contributions converge on the view that fairness outcomes depend fundamentally on the composition and coverage of training data, suggesting that governance interventions must operate not only at the level of model outputs but also at the level of data mixtures themselves.
@@ 2.3 Privacy and data protection
Privacy and data protection have long been central concerns in data-intensive systems. Early work on microdata anonymization showed that combinations of quasi-identifiers can sharply increase the risk of reidentification even when direct identifiers are removed (Samarati, 2001). Subsequent analyses of large-scale data practices demonstrate that linkage across heterogeneous sources creates cumulative disclosure risk, often in ways that are difficult to anticipate (Narayanan and Shmatikov, 2008; Rocher et al., 2019). These findings underscore that privacy loss is primarily a function of data structure, granularity, and cross-system interoperability rather than of model design alone.
A complementary line of work develops formal guarantees for limiting individual information leakage. Differential privacy provides mechanisms that bound the inference risk associated with queries or model training and quantifies how privacy loss compounds under composition (Dwork and Roth, 2014; Abowd, 2018; Wood et al., 2020). This literature makes explicit that privacy exposure depends on the magnitude, sensitivity, and reuse of the underlying data, and that even well-designed models can violate privacy when trained on highly identifying inputs.
Taken together, this literature indicates that privacy risk arises from cumulative decisions about data selection, linkage, and reuse, implying that privacy protection must often be implemented through constraints on how different data categories are combined within training systems.
@@ 2.4 Robustness and data quality
Robustness to changes in data distributions has been the focus of another active line of work. Recht et al. (2019) construct a test set for ImageNet and show that image classifiers that perform well on the original benchmark can suffer notable accuracy declines when evaluated on their new but closely related dataset. Taori et al. (2020) extend this analysis to a broader range of natural distribution shifts and document systematic performance degradation across models and training regimes. These results suggest that models can be fragile to relatively modest changes in the data generating process, especially when they have been optimized on a narrow distribution of training examples.
In parallel, research on data quality has mapped out common sources of degradation in machine learning procedures. Jakubik et al. (2024) review the state of data quality for machine learning and identify issues such as label noise, incomplete or biased coverage, missing documentation, and the inclusion of manipulated or low-veracity sources. They argue that many observed failures of deployed systems can be traced back to shortcomings in data collection and curation rather than to model architecture. These findings reinforce the view that robustness and reliability are properties of the joint data-model system, highlighting the governance importance of managing training-data composition under evolving deployment conditions.
@@ 2.5 Value of information and efficiency
A further strand of research considers the value and efficient use of information. Early work in decision analysis, notably Howard (1966), develops information value theory and shows that the expected benefit of additional information can be quantified in terms of improved decision quality. Hubbard (2014) translates these ideas into practical frameworks for measuring the value of intangibles in organizational settings, emphasizing that many institutions collect more information than they can effectively use.
In finance, Markowitz (1952) introduce the idea that portfolio choice can be framed as a balance between expected return and risk, establishing diversification and risk constraints as central features of decision under uncertainty. Subsequent research develops alternative risk measures that capture tail exposures more effectively, most notably Conditional Value-at-Risk (Rockafellar and Uryasev, 2000). Related work on coherent risk measures (Artzner et al., 1999) provides axiomatic foundations for evaluating optimal portfolio allocation. Elton et al. (2014) survey the broader development of modern portfolio theory where regulated constraints and model-based estimates jointly determine feasible allocation sets.
In the context of machine learning, classic arguments emphasize that performance gains often arise from data scale, coverage, and problem formulation rather than from incremental algorithmic refinements (Halevy et al., 2009). Ghorbani and Zou (2019) assign values to individual data points or subsets based on their marginal contribution to model performance, building on the Shapley value concept from cooperative game theory (Shapley, 1953). Subsequent work extends these ideas to federated learning, data marketplaces, and scalable privacy-aware settings under heterogeneous data (Wang et al., 2020; Sun et al., 2023; Tian et al., 2022; Wang et al., 2023). Collectively, this line of research formalizes information as an asset with heterogeneous and measurable contributions, providing a conceptual foundation for treating training data as allocatable resources subject to risk-return trade-offs.
@@ 2.6 Governance frameworks and emerging principles
Policy frameworks have begun to synthesize these research strands into articulated governance aims. The OECD Principles on Artificial Intelligence (OECD, 2019), the NIST AI Risk Management Framework (National Institute of Standards and Technology, 2023), the EU Artificial Intelligence Act (European Parliament and Council of the European Union, 2024), and the U.S. Blueprint for an AI Bill of Rights (White House Office of Science and Technology Policy, 2022) all emphasize the need for transparency and documentation of AI systems, fairness and non-discrimination in outcomes, privacy and data protection, robustness and security, and proportional use of data and computational resources. They require, to varying degrees, that high-risk AI systems document training data and model behavior, assess subgroup performance, safeguard privacy, and demonstrate robustness under realistic operating conditions.
Despite these advances, existing governance frameworks largely articulate normative objectives rather than operational mechanisms capable of sustaining them. They specify what trustworthy AI systems should achieve, but provide limited guidance on how institutions should organize, constrain, and justify the data choices underlying automated decisions. As a result, training data are frequently assembled through inherited workflows and ad hoc judgments about value and risk, with limited capacity for consistent external scrutiny. This operational gap motivates the Smart Data Portfolio framework developed in this study.
@@ 3 The Smart Data Portfolio Framework
Building on the preceding discussion, the Smart Data Portfolio framework formalizes input governance as a structured allocation problem over regulated data categories. We define a data category as a standardized class of inputs whose internal composition may change over time, but whose use is governed, constrained, and reported as a single entity. Categories can be specified by regulators at an intermediate level of detail, capturing shared provenance, governance-risk characteristics, and semantic scope, rather than raw records or model-specific features. This level of abstraction enables data inputs to be governed in a way that is stable, comparable, and auditable across sectors and institutions.
@@
SDPs with different allocations of data categories therefore generate distinct governance-performance trade-offs. As a result, some portfolios are strictly superior, in that no alternative provides higher informational return at the same level of governance-adjusted risk. The set of such portfolios for each level of risk defines the Governance-Efficient Frontier. Institutions typically select portfolios on this frontier to maximize informational return subject to the governance constraints, while regulators control the aggregate risk level.
@@
Figure 1: Illustrative governance-efficient frontier under a policy-defined risk cap. The efficient frontier (teal) traces the highest informational return attainable at each governance-risk level. The policy risk cap (blue dashed line) defines the admissible region. Filled points denote admissible SDPs; hollow points denote dominated or policy-infeasible SDPs. The governance-optimal portfolio (orange) is the highest-return point on the frontier within the risk cap.
@@
These ideas are informed by supervisory practices long established in financial regulation, where regulators define admissible risk envelopes while institutions retain discretion over internal methods. The SDP framework adapts this supervisory logic to AI inputs, formalizing how governance parameters shape feasible data allocations without constraining model choice or innovation.
4 Operationalization
This section specifies the quantities and procedures needed to implement SDPs as an auditable input-governance layer, including portfolio-level performance and governance-risk measures, the resulting governance-efficient frontier, and the regulatory constraints that define admissible data allocations.